OpenStudy (anonymous):

Write the expression in terms of sines and/or cosines and then simplify. [1+cos(x)tan(x)csc(x)]/csc(x)= ?

OpenStudy (anonymous):

\[\frac{ 1+\cos x \tan x \csc x }{ \csc x }=\frac{ 1+\cos x* \frac{ \sin x }{ \cos x }*\frac{ 1 }{ \sin x } }{ \frac{ 1 }{ \sin x }}=?\]